OpenStudy (anonymous):

-4√x+9=20 what is the answer? please i tried it but im not sure

OpenStudy (anonymous):

what answer did you get? also, is the 9 inside the radical?

OpenStudy (anonymous):

is it only x under sq rt or (x+9)...?

OpenStudy (anonymous):

it doesnt matter, both ways give complex solns

OpenStudy (anonymous):

which I am guessing is not what they are after.

OpenStudy (anonymous):

no they won't. square root of a number can be negative.

OpenStudy (anonymous):

for example -2 can be a root of 4

OpenStudy (anonymous):

lol what

OpenStudy (anonymous):

irvine, we need your feedback to help you. please respond.

OpenStudy (anonymous):

square roots cannot return negative values lol as soon as you get sqrt(anything ) = negative, theres no soln lol

OpenStudy (anonymous):

elecengineer, what is so hard to understand? \[-4\sqrt(x+9) = 20 \rightarrow \sqrt(x+9) = -5 \rightarrow x+9 = -5^2 \rightarrow x=16\]

OpenStudy (anonymous):

sub it back into the original eqn, see if it works :P

OpenStudy (anonymous):

-4(5)= 20, whoops

OpenStudy (rosey):

u should be a fan of me :)

OpenStudy (anonymous):

ah yes, this is not a function. you are right. we only consider the principal square root. my bad. ignore my answers.

OpenStudy (anonymous):

i got 'no solution' as the answer and ya the 9 is in the radical

OpenStudy (anonymous):

sorry for not anwsering previously i had to do something else